# Grothendieck’s proof of Hirzebruch-Riemann-Roch theorem

### v1

@@ 3.3 Chern and Todd classes
called the Chern classes of $\mathcal{E}$.
@@
We see that $c_{i}\left(\mathcal{E}\right)$ gives a map $A_{k}\left(X\right)\to A_{k-i}\left(X\right)$, therefore we have the feeling that the sequence of $c_{i}\left(\mathcal{E}\right)$ is bounded: it is true (namely, one can show that $c_{i}\left(\mathcal{E}\right)=0$ for $i>\mathrm{rk}\mathcal{E}$, see [Ful, Theorem 3.2, pag. 50]). Therefore the total Chern class
@@
Let $\mathcal{L}$ be an invertible sheaf on a smooth projective variety $X$, then $\mathcal{L}\simeq\mathcal{O}_{X}\left(D\right)$ for some divisor $D$. In view of Remark 3.18, we see that $c_{i}\left(\mathcal{L}\right)=0$ for $i\geq 2$, therefore
@@
Let $X=\mathbb{P}^{n}$ and let us compute $c\left(\mathcal{T}_{X}\right)$, where $\mathcal{T}$ denotes the tangent sheaf. We have the Euler exact sequence
$0\to\mathcal{O}_{X}\to\mathcal{O}_{X}\left(1\right)^{\oplus n+1}\to\mathcal{T}_{X}\to 0$
@@
$c\left(\mathcal{O}_{X}\left(1\right)^{\oplus n+1}\right)=c\left(\mathcal{O}_{X}\right)\circ c\left(\mathcal{T}_{X}\right),$
with $c\left(\mathcal{O}_{X}\right)=1$ and
$c\left(\mathcal{O}_{X}\left(1\right)^{\oplus n+1}\right)$ $=c\left(\mathcal{O}_{X}\left(1\right)^{\oplus n+1}\right)=\left(1+H\right)^{n+1}.$
@@
$c_{k}\left(\mathcal{T}_{X}\right)=\binom{n+1}{k}\cdot H^{k},$
@@
on some projective bundle over $X$. In view of Corollary 3.15, we see that the $k$-th symmetric polynomial in the $\alpha_{i}$’s (treating them as formal variables) is actually $c_{k}\left(\mathcal{E}\right)$, hence every symmetric polynomial in the $\alpha_{i}$’s in expressible in terms of the Chern classes. For a (more) formal explanation of this principle one can see [Ful, Remark 3.2.3, pag. 54]. One can also observe that, using this trick, given a morphism $f:X\to Y$ between projective varieties and a locally free coherent $\mathcal{O}_{Y}$-module $\mathcal{E}$, we have
@@
For a proof of this formulas see [Ful, Remark 3.2.3, pag. 54]
@@
as endomorphisms of $A_{\bullet}\left(X\right)\otimes\mathbb{Q}$. Ther are called the Chern and Todd character, respectively.
@@
$\mathrm{td}\left(X\right)=\mathrm{td}\left(\mathcal{T}_{X}\right)\in A_{\bullet}\left(X\right)\otimes\mathbb{Q}$
called the Todd class of $X$, where $\mathcal{T}_{X}$ is the tangent sheaf of $X$.
@@
It follows from Example 3.22 that the Chern roots of $\mathcal{T}_{\mathbb{P}^{n}}$ are just $n+1$ times the class of a hyperplane $H$, therefore
@@ 4.2 The case of smooth projective varieties
We see, for example from Proposition 2.16 and [Har, Lemma III.2.10, pag. 209], that the composition
@@
$0\to\mathcal{T}_{Y}\to j^{\ast}\mathcal{T}_{X}\to\mathcal{N}_{Y|X}\to 0$
@@
$j^{\ast}\mathrm{td}\left(X\right)=\mathrm{td}\left(j^{\ast}\mathcal{T}_{X}\right)=\mathrm{td}\left(Y\right)\cdot\mathrm{td}\left(\mathcal{N}_{Y|X}\right).$ (1)
@@
holds. In view of Proposition 2.9, we only need to check it when $y$ is the class of a locally free coherent $\mathcal{O}_{Y}$-module $\mathcal{E}$. Let
@@
which, dualized, gives181818
@@
$\sum_{k=0}^{r}\left(-1\right)^{k}\mathrm{ch}\left(\bigwedge^{k}\mathcal{L}\right)$ $=\sum_{k=0}^{r}\left(-1\right)^{k}\left(\sum_{1\leq i_{1}<\dots<i_{k}\leq n}\exp\left(a_{i_{1}}+\dots+a_{i_{k}}\right)\right)$
@@
together with the projection map $q:M\to\mathbb{P}^{1}$. For $P\neq 0$ we have $q^{-1}\left(P\right)\simeq\mathbb{P}^{n}$, while $q^{-1}\left(0\right)$ is the union of the exception divisor $E\simeq\mathbb{P}_{X}\left(\mathcal{N}_{X|\mathbb{P}^{n}}\oplus\mathcal{O}_{X}\right)$ and $Y\simeq\mathrm{Bl}_{X}\left(\mathbb{P}^{n}\right)$. Note that the product $j\times\mathbb{P}^{1}:X\times\mathbb{P}^{1}\to\mathbb{P}^{n}\times\mathbb{P}^{1}$ makes the preimage of $X\times\left\{0\right\}$ locally principal, hence there exists a unique $v:X\times\mathbb{P}^{1}\to M$ making the diagram
@@
$X\times\left\{0\right\}\simeq X\to\mathbb{P}_{X}\left(\mathcal{N}_{X|\mathbb{P}^{n}}\oplus\mathcal{O}_{X}\right)$
@@
hence we have a relation
$\sum_{k=0}^{r}\left(-1\right)^{k}\mathrm{ch}\left(\mathcal{F}_{k|Y}\right)\cdot\left[Y\right]+\sum_{k=0}^{r}\left(-1\right)^{k}\mathrm{ch}\left(\mathcal{F}_{k|E}\right)\cdot\left[E\right]=\sum_{k=0}^{r}\left(-1\right)^{k}\mathrm{ch}\left(\mathcal{F}_{k|\mathbb{P}^{n}}\right)\cdot\left[\mathbb{P}^{n}\right].$
@@
is exact and hence $\sum_{k=0}^{r}\left(-1\right)^{k}\mathrm{ch}\left(\mathcal{F}_{k|Y}\right)\cdot\left[Y\right]=0$. In view of Lemma 4.5 we have
$\sum_{k=0}^{r}\left(-1\right)^{k}\mathrm{ch}\left(\mathcal{F}_{k|E}\right)\cdot\left[E\right]=\frac{\mathrm{ch}\left(\mathcal{E}\right)}{\mathrm{td}\left(\mathcal{N}_{X|\mathbb{P}^{n}}\right)}\cdot\left[X\right],$
therefore
$\frac{\mathrm{ch}\left(\mathcal{E}\right)}{\mathrm{td}\left(\mathcal{N}_{X|\mathbb{P}^{n}}\right)}\cdot\left[X\right]=\sum_{k=0}^{r}\left(-1\right)^{k}\mathrm{ch}\left(\mathcal{F}_{k|\mathbb{P}^{n}}\right)\cdot\left[\mathbb{P}^{n}\right]$
which, pushed forward along $g:M\to\mathbb{P}^{n}$, gives the relation as in the hypothesis of Lemma 4.4. ∎
@@ 4.3.1 Explicit formulas in low dimension
Theorem 4.7 ((Riemann-Roch for curves)).
Let $\mathcal{L}\in\mathrm{Pic}\left(C\right)$ be locally free sheaf on $C$ of rank 1, then
$\chi\left(C,\mathcal{L}\right)=\deg\mathcal{L}+\chi\left(C,\mathcal{O}_{C}\right).$
@@
We only need to see that
$\int_{C}\mathrm{ch}\left(\mathcal{L}\right)\cdot\mathrm{td}\left(\mathcal{C}\right)=\deg\mathcal{L}+\chi\left(C,\mathcal{O}_{C}\right).$
Let $D$ be a divisor such that $\mathcal{L}\simeq\mathcal{O}_{C}\left(D\right)$, then
$\mathrm{ch}\left(\mathcal{L}\right)=1+c_{1}\left(\mathcal{L}\right)$
and
$\mathrm{td}\left(C\right)=1+\frac{1}{2}c_{1}\left(\mathcal{T}_{C}\right).$
Since $C$ has dimension 1, we have $A_{2}\left(C\right)=0$ and hence
$\int_{C}\mathrm{ch}\left(\mathcal{L}\right)\cdot\mathrm{td}\left(C\right)$ $=\int_{C}\left(1+c_{1}\left(\mathcal{L}\right)\right)\cdot\left(1+\frac{1}{2}c_{1}\left(\mathcal{T}_{C}\right)\right)$
$=\int_{C}\left(1+c_{1}\left(\mathcal{L}\right)+\frac{1}{2}c_{1}\left(\mathcal{T}_{C}\right)\right)$
$=\int_{C}\left(1+c_{1}\left(\mathcal{L}\right)-\frac{1}{2}c_{1}\left(\Omega_{C}^{1}\right)\right)$
$=\int_{C}\left(c_{1}\left(\mathcal{L}\right)-\frac{1}{2}c_{1}\left(\Omega_{C}^{1}\right)\right)$
$=\deg\left(D\right)-\frac{1}{2}\left(2g-2\right)$
$=\deg\left(D\right)-g+1.$
Theorem 4.8 ((Riemann-Roch for surfaces)).
@@
$\chi\left(S,\mathcal{L}\right)=\frac{1}{2}\mathcal{L}\cdot\left(\mathcal{L}\otimes_{\mathcal{O}_{S}}\omega_{S}^{\lor}\right)+\frac{1}{12}\left(K_{S}^{2}+\int_{S}c_{2}\left(\mathcal{T}_{S}\right)\right).$
In particular
$\chi\left(S,\mathcal{O}_{S}\right)=\frac{1}{12}\left(K_{S}^{2}+\int_{S}c_{2}\left(\mathcal{T}_{S}\right)\right).$
@@
$\int_{S}\mathrm{ch}\left(\mathcal{L}\right)\cdot\mathrm{td}\left(S\right)$ $=\int_{S}\left(1+c_{1}\left(\mathcal{L}\right)+\frac{1}{2}c_{1}\left(\mathcal{L}\right)^{2}\right)\cdot\left(1+\frac{1}{2}c_{1}\left(\mathcal{T}_{S}\right)+\frac{1}{12}\left(c_{1}\left(\mathcal{T}_{S}\right)^{2}+c_{2}\left(\mathcal{T}_{S}\right)\right)\right)$
$=\int_{S}\left(1+\frac{1}{2}c_{1}\left(\mathcal{L}\right)^{2}+\frac{1}{2}c_{1}\left(\mathcal{L}\right)\cdot c_{1}\left(\mathcal{T}_{S}\right)+\frac{1}{12}\left(c_{1}\left(\mathcal{T}_{S}\right)^{2}+c_{2}\left(\mathcal{T}_{S}\right)\right)\right)$
$=\frac{1}{2}D^{2}-\frac{1}{2}D\cdot K_{S}+\frac{1}{12}\left(K_{S}^{2}+\int_{S}c_{2}\left(\mathcal{T}_{S}\right)\right).$

### v2

@@ 3.4 Chern and Todd classes
called the Chern classes of $\mathcal{E}$ [^12] You may wonder why such a complicated recursive definition. One could define the total Segre class $s\left(\mathcal{E}\right)=\sum_{i\geq 0}s_{i}\left(\mathcal{E}\right)$ that turns out to be an isomorphism. Its inverse is the total Chern class and it’s given by $c\left(\mathcal{E}\right)=\sum_{i\geq 0}c_{i}\left(\mathcal{E}\right)$ of Remark 3.18..
@@
We see that $c_{i}\left(\mathcal{E}\right)$ gives a map $A_{k}\left(X\right)\to A_{k-i}\left(X\right)$, therefore we have the feeling that the sequence of $c_{i}\left(\mathcal{E}\right)$ is bounded: it is true (namely, one can show that $c_{i}\left(\mathcal{E}\right)=0$ for $i>\mathrm{rk}\left(\mathcal{E}\right)$, see [Ful, Theorem 3.2, pag. 50]). Therefore the total Chern class
@@
Let $\mathcal{L}$ be an invertible sheaf on $X$, then $\mathcal{L}\simeq\mathcal{O}_{X}\left(D\right)$ for some divisor $D$. In view of Remark 3.18, we see that $c_{i}\left(\mathcal{L}\right)=0$ for $i\geq 2$, therefore
@@
Let us compute $c\left(\mathrm{T}_{\mathbb{P}^{n}}\right)$, where $\mathrm{T}$ denotes the tangent sheaf. We have the Euler exact sequence
$0\to\mathcal{O}_{\mathbb{P}^{n}}\to\mathcal{O}_{\mathbb{P}^{n}}\left(1\right)^{\oplus n+1}\to\mathrm{T}_{\mathbb{P}^{n}}\to 0$
@@
$c\left(\mathcal{O}_{\mathbb{P}^{n}}\left(1\right)^{\oplus n+1}\right)=c\left(\mathcal{O}_{\mathbb{P}^{n}}\right)\circ c\left(\mathrm{T}_{\mathbb{P}^{n}}\right),$
with $c\left(\mathcal{O}_{\mathbb{P}^{n}}\right)=1$ and
$c\left(\mathcal{O}_{\mathbb{P}^{n}}\left(1\right)^{\oplus n+1}\right)$ $=\left(1+H\right)^{n+1}.$
@@
$c_{k}\left(\mathrm{T}_{\mathbb{P}^{n}}\right)=\binom{n+1}{k}\cdot H^{k},$
@@
on some projective bundle over $X$. In view of Corollary 3.15, we see that the $k$-th symmetric polynomial in the $\alpha_{i}$’s (treating them as formal variables) is actually $c_{k}\left(\mathcal{E}\right)$, hence every symmetric polynomial in the $\alpha_{i}$’s in expressible in terms of the Chern classes. For a (more) formal explanation of this principle one can see [Ful, Remark 3.2.3, pag. 54]. We also point out that, using this trick, given a morphism $f:X\to Y$ between projective varieties and a locally free coherent $\mathcal{O}_{Y}$-module $\mathcal{E}$, we have
@@
In particular $\det\left(\mathcal{E}\right)$ has Chern root $\alpha_{1}+\dots+\alpha_{n}$. For a proof of this formulas see [Ful, Remark 3.2.3, pag. 54]
@@
as endomorphisms of $A_{\bullet}\left(X\right)\otimes\mathbb{Q}$. They are called the Chern and Todd character, respectively.
@@
Now that we view Chern classes as elements of the Chow ring $A_{\bullet}(X)$, we can define the degree of a locally free sheaf $\mathcal{E}$ as
$\deg\left(\mathcal{E}\right)=\deg\left(c_{1}\left(\mathcal{E}\right)\right).$
Using the splitting principle (Remark 3.23 and Example 3.24), one can verify that
$c_{1}(\mathcal{E})=c_{1}(\det\mathcal{E}).$
Thus, this definition coincides with the classical definition of the degree as the degree of the determinant line bundle.
Remark 3.29.
@@
$\mathrm{td}\left(X\right)=\mathrm{td}\left(\mathrm{T}_{X}\right)\in A_{\bullet}\left(X\right)\otimes\mathbb{Q}$
called the Todd class of $X$, where $\mathrm{T}_{X}$ is the tangent sheaf of $X$.
@@
In view of the multiplicativity of the Todd class (Proposition 3.27) applied to the Euler exact sequence (from Example 3.22) we get:
$\mathrm{td}(\mathcal{O}_{\mathbb{P}^{n}}\left(1\right)^{\oplus n+1})=\mathrm{td}(\mathcal{O}_{\mathbb{P}^{n}})\cdot\mathrm{td}(\mathrm{T}_{\mathbb{P}^{n}})$
The Todd class of the trivial bundle $\mathrm{td}(\mathcal{O}_{\mathbb{P}^{n}})$ is $1$. For the sum, we use multiplicativity (Proposition 3.27) and the definition of the Todd class for a line bundle (Definition 3.25) whose first Chern class is $H$:
$\mathrm{td}(\mathcal{O}_{\mathbb{P}^{n}}\left(1\right)^{\oplus n+1})=\mathrm{td}(\mathcal{O}_{\mathbb{P}^{n}}(1))^{n+1}=\left(\frac{H}{1-\exp(-H)}\right)^{n+1}$
Substituting this back, we obtain the formula:
$\mathrm{td}\left(\mathrm{T}_{\mathbb{P}^{n}}\right)=\frac{H^{n+1}}{\left(1-\exp\left(-H\right)\right)^{n+1}},$
@@ 4.2 The case of smooth projective varieties
We see, in view of Proposition 2.17, that the composition
@@
$0\to\mathrm{T}_{Y}\to j^{\ast}\mathrm{T}_{X}\to\mathcal{N}_{Y|X}\to 0$
@@
$j^{\ast}\mathrm{td}\left(X\right)=\mathrm{td}\left(j^{\ast}\mathrm{T}_{X}\right)=\mathrm{td}\left(Y\right)\cdot\mathrm{td}\left(\mathcal{N}_{Y|X}\right).$ (1)
@@
holds. In view of Theorem 2.8, we only need to check it when $y$ is the class of a locally free coherent $\mathcal{O}_{Y}$-module $\mathcal{E}$. Let
@@
which, dualized, gives171717
@@
$\sum_{k=0}^{r}\left(-1\right)^{k}\mathrm{ch}\left(\bigwedge^{k}\mathcal{L}\right)$ $=\sum_{k=0}^{r}\left(-1\right)^{k}\left(\sum_{1\leq i_{1}<\dots<i_{k}\leq r}\exp\left(a_{i_{1}}+\dots+a_{i_{k}}\right)\right)$
@@
together with the projection map $q:M\to\mathbb{P}^{1}$. For $P\neq 0$ we have $q^{-1}\left(P\right)\simeq\mathbb{P}^{n}$, while $q^{-1}\left(0\right)$ is the union of the exception divisor $E\simeq\mathbb{P}_{X}\left(\mathcal{N}_{X|\mathbb{P}^{n}}\oplus\mathcal{O}_{X}\right)$ and $Y\simeq\mathrm{Bl}_{X}\left(\mathbb{P}^{n}\right)$.
Let $i_{E}:E\to M$, $i_{Y}:Y\to M$ be the inclusions of the components of the fiber over 0, and let $i_{\infty}:\mathbb{P}^{n}\to M$ be the inclusion of the fiber over $\infty$.
Note that the product $j\times\mathbb{P}^{1}:X\times\mathbb{P}^{1}\to\mathbb{P}^{n}\times\mathbb{P}^{1}$ makes the preimage of $X\times\left\{0\right\}$ locally principal, hence there exists a unique $v:X\times\mathbb{P}^{1}\to M$ making the diagram
@@
$v_{0}:X\simeq X\times\left\{0\right\}\to E\simeq\mathbb{P}_{X}\left(\mathcal{N}_{X|\mathbb{P}^{n}}\oplus\mathcal{O}_{X}\right)$
@@
where $[E]$, $[Y]$, $\mathbb{P}^{n}]$ denote the classes of the images of $i_{E}$, $_{Y}, ii_{\infty}$ in $A_{\bullet}(M)$. This gives the relation in $A_{\bullet}(M)\otimes\mathbb{Q}$:
$i_{Y*}\left(\sum_{k=0}^{r}\left(-1\right)^{k}\mathrm{ch}\left(\mathcal{F}_{k|Y}\right)\right)+i_{E*}\left(\sum_{k=0}^{r}\left(-1\right)^{k}\mathrm{ch}\left(\mathcal{F}_{k|E}\right)\right)=i_{\infty*}\left(\sum_{k=0}^{r}\left(-1\right)^{k}\mathrm{ch}\left(\mathcal{F}_{k|\mathbb{P}^{n}}\right)\right).$
@@
is exact and hence $i_{Y*}\left(\sum_{k=0}^{r}\left(-1\right)^{k}\mathrm{ch}\left(\mathcal{F}_{k|Y}\right)\right)=0$. The term for $E$ is $i_{E*}\left(\mathrm{ch}(v_{0!}(\mathcal{E}))\right)$, since $\mathcal{F}_{\bullet|E}$ is a resolution of $v_{0*}\mathcal{E}$. By Lemma 4.5, this is $i_{E*}\left(v_{0*}\left(\dfrac{\mathrm{ch}(\mathcal{E})}{\mathrm{td}(\mathcal{N}_{X|E})}\right)\right)$, where $\mathcal{N}_{X|E}\simeq\mathcal{N}_{X|\mathbb{P}^{n}}$. The term for $\mathbb{P}^{n}$ is $i_{\infty*}\left(\mathrm{ch}(j_{!}(\mathcal{E}))\right)$, since $\mathcal{F}_{\bullet|\mathbb{P}^{n}}$ is a resolution of $v_{\infty*}\mathcal{E}\simeq j_{*}\mathcal{E}$.
The relation in $A_{\bullet}(M)\otimes\mathbb{Q}$ thus simplifies to:
$(i_{E}\circ v_{0})_{*}\left(\frac{\mathrm{ch}\left(\mathcal{E}\right)}{\mathrm{td}\left(\mathcal{N}_{X|\mathbb{P}^{n}}\right)}\right)=i_{\infty*}\left(\mathrm{ch}(j_{!}\mathcal{E})\right)$
Let $g:M\to\mathbb{P}^{n}$ be the projection map (il blow-down). We apply the pushforward $g_{*}$ to both sides of the equation. By functoriality of the pushforward (Lemma 3.5) and noting that $g\circ i_{E}\circ v_{0}=j$ (the original embedding) and $g\circ i_{\infty}=\mathrm{id}_{\mathbb{P}^{n}}$, we get:
$j_{*}\left(\frac{\mathrm{ch}\left(\mathcal{E}\right)}{\mathrm{td}\left(\mathcal{N}_{X|\mathbb{P}^{n}}\right)}\right)=\mathrm{ch}(j_{!}\mathcal{E})$
which, by Proposition 3.27 applied to the resolution $\mathcal{F}_{\bullet|\mathbb{P}^{n}}$, is exactly the relation
$j_{*}\left(\frac{\mathrm{ch}\left(\mathcal{E}\right)}{\mathrm{td}\left(\mathcal{N}_{X|\mathbb{P}^{n}}\right)}\right)=\sum_{k=0}^{r}\left(-1\right)^{k}\mathrm{ch}\left(\mathcal{F}_{k|\mathbb{P}^{n}}\right)$
required in the hypothesis of Lemma 4.4. ∎
@@ 4.3.1 Explicit formulas in low dimension
Theorem 4.7 (Riemann-Roch for curves).
Let $\mathcal{E}$ be locally free sheaf on $C$ of rank r, then
$\chi(C,\mathcal{E})=\deg(\mathcal{E})+r\cdot(1-g)$
@@
In view of Theorem 4.6 we need to see that
$\int_{C}\mathrm{ch}(\mathcal{E})\cdot\mathrm{td}(\mathrm{T}_{C})=\deg(\mathcal{E})+r\cdot(1-g)$
We have $A_{\bullet}(C)=A_{1}(C)\oplus A_{0}(C)$ with $A_{0}(C)\cdot A_{0}(C)=0$. In view of Definition 3.25 we have
$\mathrm{td}(\mathrm{T}_{C})=1+\frac{1}{2}c_{1}(\mathrm{T}_{C}),$
Moreover since $\mathrm{T}_{C}=(\omega_{C})^{\lor}$, we see that $\deg(c_{1}(\mathrm{T}_{C}))=-\deg(\omega_{C})=2-2g$. Again in view of Definition 3.25 $\mathrm{ch}(\mathcal{E})=\mathrm{rank}(\mathcal{E})+c_{1}(\mathcal{E})=r+c_{1}(\mathcal{E})$, whence
$\mathrm{ch}(\mathcal{E})\cdot\mathrm{td}(\mathrm{T}_{C})=(r+c_{1}(\mathcal{E}))\cdot\left(1+\frac{1}{2}c_{1}(\mathrm{T}_{C})\right)=r+c_{1}(\mathcal{E})+\frac{r}{2}c_{1}(\mathrm{T}_{C})$
since $c_{1}(\mathcal{E})\cdot\dfrac{1}{2}c_{1}(\mathrm{T}_{C})\in A_{-1}(C)=0$. Finally we compute
$\int_{C}\mathrm{ch}(\mathcal{E})\cdot\mathrm{td}(\mathrm{T}_{C})$ $=\deg\left(c_{1}(\mathcal{E})+\frac{r}{2}c_{1}(\mathrm{T}_{C})\right)$
$=\deg(c_{1}(\mathcal{E}))+\frac{r}{2}\deg(c_{1}(\mathrm{T}_{C}))$
$=\deg(\mathcal{E})+r(1-g).$
Theorem 4.8 (Riemann-Roch for surfaces).
@@
$\chi\left(S,\mathcal{L}\right)=\frac{1}{2}\mathcal{L}\cdot\left(\mathcal{L}\otimes_{\mathcal{O}_{S}}\omega_{S}^{\lor}\right)+\chi\left(S,\mathcal{O}_{S}\right),$
moreover
$\chi\left(S,\mathcal{O}_{S}\right)=\frac{1}{12}\left(K_{S}^{2}+\int_{S}c_{2}\left(\mathrm{T}_{S}\right)\right).$
@@
$\int_{S}\mathrm{ch}\left(\mathcal{L}\right)\cdot\mathrm{td}\left(S\right)$ $=\int_{S}\left(1+c_{1}\left(\mathcal{L}\right)+\frac{1}{2}c_{1}\left(\mathcal{L}\right)^{2}\right)\cdot\left(1+\frac{1}{2}c_{1}\left(\mathrm{T}_{S}\right)+\frac{1}{12}\left(c_{1}\left(\mathrm{T}_{S}\right)^{2}+c_{2}\left(\mathrm{T}_{S}\right)\right)\right)$
$=\int_{S}\left(1+\frac{1}{2}c_{1}\left(\mathcal{L}\right)^{2}+\frac{1}{2}c_{1}\left(\mathcal{L}\right)\cdot c_{1}\left(\mathrm{T}_{S}\right)+\frac{1}{12}\left(c_{1}\left(\mathrm{T}_{S}\right)^{2}+c_{2}\left(\mathrm{T}_{S}\right)\right)\right)$
$=\frac{1}{2}D^{2}-\frac{1}{2}D\cdot K_{S}+\frac{1}{12}\left(K_{S}^{2}+\int_{S}c_{2}\left(\mathrm{T}_{S}\right)\right).$